OpenStudy (anonymous):

How do I factor this expression? 6x^2+7x-20. I know the answer is (3x - 4)(2x + 5) because I FOILed it, but otherwise, what would I have to do?

OpenStudy (anonymous):

the easiest method will be to use the quadratic formula and then work backwards

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

its a pleasure..happy 2 help

OpenStudy (anonymous):

How would I factor it after using the quadratic formula? I got -5/2 and 4/3.

OpenStudy (anonymous):

then you would work backwards and say one of the factors is \(x+\frac{5}{2}\) but instead multiply by 2 and say it is \(2x+5\)

OpenStudy (anonymous):

similarly for \(\frac{4}{3}\) you can write \[x=\frac{4}{3}\iff x-\frac{4}{3}=0\iff 3x-4=0\]

OpenStudy (anonymous):

Thank you! I understand now! (:

OpenStudy (anonymous):

sometimes it is easier just to factor, but the quadratic formula always works

OpenStudy (anonymous):

yw

OpenStudy (campbell_st):

well here is an easy method that works everytime for a quadratic \[ax^2 + bx + c \] start my multiplying a and c in your question its 6 x -20 = -120 find the factors of -120 that add to 7 the larger factor is positive the smaller negative so its 15 and - 8 next write (ax + factor 1)(ax + factor 2) -------------------------- a so for your question you have (6x + 15)(6x - 8) ---------------- 6 remove common factors from the binomials 3(2x + 5)2(3x - 4) 3*2(2x + 5)(3x -4) ----------------- = ------------------ 6 6 cancel the common factor leaves (2x + 5)(3x -4) no guessing and no GQF needed.